Solve the system of equations by elimination. The answer needs to be in the ordered pair of (x,y,z) form.
x + y + z = 6
x + 2y - z = 1
3x - 3y - z = -32


Answer by lwsshak3(11628)

Solve the system of equations by elimination. The answer needs to be in the ordered pair of (x,y,z) form.
x + y + z = 6 (eq1)
x + 2y - z = 1 (eq2
3x - 3y - z = -32 (eq)3
***
x + y + z = 6 (eq1)
x + 2y - z = 1 (eq2)
add to eliminate z
2x+3y=7
..
x + 2y - z = 1 (eq2)
3x - 3y - z = -32 (eq3)
subtract to eliminate z
-2x+5y=33
..
-2x+5y=33
2x+3y=7
add to eliminate x
8y=40
y=5
2x=7-3y=7-15=-8
x=-4
z=6-x-y=6+4-5=5
solution:(x,y,z)=(-4,5,5)